# Controlling the quantum phase transition in a double quantum dot Josephson junction via interactions

### v1

@@ Abstract
In this work, we employ a surrogate BCS model with discrete energy levels to investigate a hybrid system comprising two quantum dots (QD1 and QD2), where QD1 is tunnel-coupled to two superconducting leads. Through exact diagonalization of this system, we obtain numerically exact solutions that enable rigorous computation of key physical quantities. Our analysis reveals a rich phase diagram featuring multiple controllable phase transitions mediated by quantum dot interactions. Specifically, the system first undergoes an initial phase transition when tuning QD2’s interaction strength while maintaining QD1 in the non-interacting regime. Subsequent adjustment of QD1’s interaction induces a secondary phase transition, followed by a third transition arising from inter-dot coupling modulation. Furthermore, we demonstrate that parallel magnetic field application can drive reversible ferromagnetic-antiferromagnetic phase transitions under specific parameter conditions. Finally, we report the emergence of non-local magnetization phenomena when subjecting QD1 to weak magnetic fields. And our results demonstrate that the orientation of nonlocal magnetization can be precisely manipulated through systematic adjustment of the on-site interaction strength $U_{2}$ in QD2.
@@ I Introduction
The Josephson junction represents a paradigmatic mesoscopic system comprising a central weak-link region sandwiched between two superconducting electrodes. This fundamental structure has served as a versatile platform for investigating coherent quantum transport phenomena, as evidenced by extensive theoretical [1, 2, 3, 4, 5, 6, 7, 8, 9] and experimental [10, 11, 12, 13, 14]studies. A hallmark characteristic of Josephson junctions is the generation of a dissipationless supercurrent—the Josephson current—that flows without any applied voltage bias when a phase difference exists between the superconducting terminals[15, 16]. This equilibrium current arises from phase-coherent Andreev reflection processes at the superconductor-normal interfaces. A particularly intriguing phenomenon in these systems is the $0-\pi$ quantum phase transition [17, 18, 19, 20, 21, 22, 23, 24, 25, 26], which manifests as a sign reversal of the Josephson current and has been unambiguously observed in experiments [27, 28, 29, 30, 31, 32, 33]. This transition fundamentally reflects a change in the ground state parity of the system. In quantum dot(QD)-based Josephson junctions, two distinct control parameters can drive this transition: (i) tuning the on-dot Coulomb interaction strength [17, 18, 19, 20, 21, 22], or (ii) applying and varying an external magnetic field [23, 24, 25, 26]. Remarkably, despite their different control mechanisms, both approaches share a common microscopic origin—they induce a quantum phase transition(QPT) between two distinct ground state configurations: a singlet state (formed by Cooper pairs via Andreev bound states) and a doublet state (characterized by an unpaired electron). This transition can be understood within a unified framework of competition between Kondo screening or magnetic field and superconducting correlations, where the relative strength of these effects determines the ground state parity.
@@
While QPTs in Josephson junctions containing a single QD in their central region are now well understood, significant research attention has shifted to double-QD systems, as evidenced by growing theoretical [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51] and experimental[52, 53, 54, 55, 56, 57] investigations. The enhanced complexity of ground state physics in these two-QD systems gives rise to QPT phenomena that extend beyond conventional $0-\pi$ transitions, manifesting in three primary configurations: (i) the series configuration [34, 35, 36, 37, 38, 39, 40], where interconnected QDs separately link to each superconducting lead; (ii) the parallel configuration [41, 42, 43, 44, 45, 46], featuring both QDs jointly connected to each superconducting lead; and (iii) the Fano-type configuration [47, 48, 49], where QD1 bridges the superconductors while QD2 remains indirectly coupled. Recent work on Fano-type double interacting QD josephson junction [48] has revealed both a second-order superconducting proximity effect and QPTs induced through QD2’s electron interaction tuning, yet crucially omitted examination of QD1’s on-site interactions and inter-dot interactions between QD1 and QD2 effects - either of which could drive additional QPTs and represent compelling research avenues. Furthermore, given established magnetic-field-induced QPTs in single-QD junctions, the analogous response in double-QD systems remains an open question of considerable fundamental interest, presenting an important unexplored dimension in hybrid superconducting nanostructure research.
@@
Historically, theoretical analysis of Josephson junctions incorporating interacting QDs posed significant challenges due to the inherent complexity of treating electron correlations. Until recently, the typically effective approaches were Wilson’s numerical renormalization group (NRG) [58, 59, 48, 42] and quantum Monte Carlo (QMC) [60, 61, 62] methods, both of which are computationally demanding. The emergence of the surrogate model technique [63, 64, 65, 66, 67, 68] has provided a simpler yet rigorous alternative to these conventional methods. This innovative approach transforms the continuous leads Hamiltonian into a discrete analog through systematic discretization, enabling exact diagonalization of the complete system Hamiltonian. The resulting eigenvalues and eigenstates provide direct access to all relevant physical quantities through straightforward computation. Importantly, this method has already demonstrated its effectiveness for single-QD Josephson junctions, and its fundamental premise - the faithful representation of superconducting leads through equivalent discrete Hamiltonians - strongly suggests equal validity for double-QD systems. In the present study, we implement this approach by discretizing each superconducting lead into three equivalent sites, then employ state-space expansion techniques to perform exact diagonalization of the full Hamiltonian for a Fano-coupled double-QD Josephson junction. This methodology allows comprehensive computation of system observables, which we subsequently analyze to investigate: (i) interaction-driven QPTs under various quantum dot coupling regimes, and (ii) magnetic-field-induced phase transition phenomena in this novel nanoscale system."
@@ II Model Hamiltonian and Theoretical Method
$\begin{split}H_{QD}=&\sum_{i\sigma}\varepsilon_{i\sigma}d^{{\dagger}}_{i\sigma}d_{i\sigma}+\sum_{i}U_{i}n_{i\uparrow}n_{i\downarrow}\\ &+U_{12}n_{1}n_{2}+t\sum_{\sigma}d^{{\dagger}}_{1\sigma}d_{2\sigma}+h.c.,\end{split}$ (2)
@@
The operator $d^{{\dagger}}_{i\sigma}$($d_{i\sigma}$) creates (annihilates) an electron with spin $\sigma$ and energy $\varepsilon_{i\sigma}$ on the i-th quantum dot (QD), where $U_{i}$ represents the on-site Coulomb repulsion. The interdot tunneling is characterized by the hopping amplitude $t$, while $U_{12}$ denotes the interdot Coulomb interaction, which is typically weaker than the intradot interactions($U_{12}<U_{i}$). Similarly, $c_{\eta\bm{k}\sigma}^{{\dagger}}$($c_{\eta\bm{k}\sigma}$) creates (annihilates) an electron with spin $\sigma$, momentum $\bm{k}$, and energy $\varepsilon_{\eta\bm{k}\sigma}$ in the $\eta$-th SC lead, where $\Delta e^{i\phi_{\eta}}$ represents the complex superconducting order parameter, where $\Delta$ and $\phi$ are real numbers. The SC-QD coupling is described by the tunneling Hamiltonian $H_{T}$ with momentum-independent tunneling matrix elements $V$. We approximate the momentum summation by an energy integral using a constant density of states $\rho_{F}=\frac{1}{2D}$, where $D$ is the half-bandwidth. Throughout this work, we employ natural units with $\hbar$ = $k_{B}$ = $e$ = 1.
@@
the $\tilde{g}$ function is obtained by integrating out an effective superconducting bath that has the same gap $\Delta$ as the original system. This bath consists of three discrete energy levels, located at $\xi_{0}$ and ±$\pm|\xi_{1}|$, which are coupled to the dot via a tunneling matrix element $\tilde{t}_{l}=\sqrt{\gamma_{1}\Gamma}/2$. Since the number of discrete levels is odd, it follows that $\xi_{0}=0$ [68]. The effective bath is thus characterized by the parameters $\{{\tilde{\gamma}_{0}},{\tilde{\gamma}_{1}},\tilde{\xi}_{0},\tilde{\xi}_{1}\}$, which are determined by minimizing the cost function:
@@
Once a satisfactory fit is obtained, the parameters $\{{\tilde{\gamma}_{0}},{\tilde{\gamma}_{1}},\tilde{\xi}_{0},\tilde{\xi}_{1}\}$ define a surrogate model Hamiltonian, representing a discretized version of the original system. This is achieved by replacing the continuous momentum variable $\bm{k}$ with a discrete index $l$, such that $\xi_{\bm{k}}\rightarrow\xi_{l}$ and $t\rightarrow\tilde{t}_{l}$. The resulting Hamiltonian takes the form:
@@ III Results and Discussions
In this work, we numerically solve the low-dimensional surrogate model through exact diagonalization and subsequently calculate the system’s thermodynamic properties including entropy, parity, spin correlations, QD pair term, QD occupation number, and Josephson current, with particular emphasis on the particle-hole symmetric point where $\varepsilon_{i}=-\frac{U_{i}}{2}$. Following standard conventions, we establish $\phi_{L}=\phi$ and $\phi_{R}=0$ for the superconducting phases, with $\phi=0$ as the default value, and maintain the system at zero temperature unless otherwise specified. Furthermore, we fix the tunneling rate at $\Gamma=2$ and adopt the superconducting gap $\Delta=1$ as our fundamental energy scale unless specifically stated.
@@ III.1 Single QD josephson junction
$I_{J}=\frac{\partial E_{S}}{\partial\phi},$ (15)
@@
III.2 QPT are induced by varying the interaction parameters
For the subsequent calculations, we fix the interdot tunneling parameter at $t=1$ and examine the system’s physical properties when the interaction strength in QD1 is zero ($U_{1}=0$).
@@ III.2 QPT are induced by varying the interaction parameters
Fig. 3: (Colour online) System properties analysis: (a) entropy at fixed temperature $T=0.005$, (b) system parity operator expectation value $P$ for various magnetic field strength $B$, (c) spin correlation function $\langle\vec{S}_{1}\!\!\cdot\!\!\vec{S}_{2}\rangle$, (d) normalized superconducting order parameters, $\langle d_{1\downarrow}d_{1\uparrow}\rangle/\Delta$ for QD1 and $\langle d_{2\downarrow}d_{2\uparrow}\rangle/\Delta$ for QD2 and (e) Josephson current $I_{J}/I_{0}$ (where $I_{0}\equiv e\Delta/\hbar$) at fixed superconducting phase difference $\phi=0.3\pi$ as a function of QD2 interaction strength $U_{2}/\Delta$. (f) $I_{J}/I_{0}$ versus phase difference $\phi$ for varying QD2 interaction strengths. The remaining system parameters are fixed at interdot tunneling $t=1$, on-site interaction $U_{1}=0$, and interdot interaction $U_{12}=0$.
@@
Figure 3(a) displays the entropy as a function of QD2 interaction strength $U_{2}$ at fixed temperature $T=0.005$. Throughout Fig. 3, we maintain $U_{1}=0$ for QD1 and $U_{12}=0$ for interdot Coulomb interaction. The entropy exhibits a QPT near $U_{2}\approx 5\Delta$, with three distinct regimes: (i) $S=0$ (singlet state) for $U_{2}<5\Delta$, (ii) $S=\ln 3$ at the critical point, and (iii) $S=\ln 2$ (doublet state) for $U_{2}>5\Delta$. The enhanced entropy value of $\ln 3$ at the quantum critical point, as previously discussed, arises from the emergent threefold degeneracy in the ground state manifold during the phase transition. Figure 3(b) shows the parity $P$ versus $U_{2}$. Post-transition, the system enters a doubly degenerate doublet state. To enable parity measurement, we apply a weak magnetic field to QD1, lifting the degeneracy. The parity transitions sharply from 00 (even) to ±$\pm 1$ (odd) at the QPT, confirming the change in ground state. The interdot spin correlation in Fig. 3(c) plots an antiferromagnetic configuration ($\langle\vec{S}_{1}\!\cdot\!\vec{S}_{2}\rangle<0$) before the QPT that vanishes ($\langle\vec{S}_{1}\!\cdot\!\vec{S}_{2}\rangle\approx 0$) in the doublet phase. The spin correlation behavior emerges from distinct ground state configurations in different phases: (i) In the singlet phase prior to the QPT, the formation of spatially localized Cooper pairs between QD1 and QD2 yields antiferromagnetic spin alignment; (ii) In the post-transition doublet phase where QD1 enters an ABS, the spin correlations become negligible. Illustrated in Fig. 3(d) is the normalized superconducting order parameters $\langle d_{1\downarrow}d_{1\uparrow}\rangle/\Delta$ (QD1) and $\langle d_{2\downarrow}d_{2\uparrow}\rangle/\Delta$ (QD2) versus $U_{2}$. At weak interactions, both orders are significant, but they decrease with increasing $U_{2}$ due to Coulomb suppression of double occupancy and empty occupancy. After the QPT, $\langle d_{1\downarrow}d_{1\uparrow}\rangle$ surges while $\langle d_{2\downarrow}d_{2\uparrow}\rangle$ vanishes, reflecting the system’s transition to a state where QD2 reaches single occupancy (particle-hole symmetry point) while QD1 forms Andreev bound states.
@@
Through rigorous analysis of the aforementioned physical quantities, we establish that the system’s ground state prior to the QPT or in the absence of electronic correlations within the QDs can be precisely described by the superposition state: $A_{1}|0_{1}0_{2}\rangle+B_{1}|0_{1}d_{2}\rangle+C_{1}|d_{1}0_{2}\rangle+D_{1}|d_{1}d_{2}\rangle+E_{1}(|\!\!\uparrow_{1}\downarrow_{2}\rangle-|\!\!\downarrow_{1}\uparrow_{2}\rangle)$, where the coefficients $A_{1}$, $B_{1}$, $C_{1}$, $D_{1}$, and $E_{1}$ quantitatively characterize the relative contributions of each entangled quantum state. Remarkably, the system undergoes a dramatic reorganization during the QPT, with the ground state abruptly collapsing into the simplified forms $|0_{1}\!\!\uparrow_{2}\rangle-|d_{1}\!\!\uparrow_{2}\rangle$ and $|0_{1}\!\!\downarrow_{2}\rangle-|d_{1}\!\!\downarrow_{2}\rangle$. This transition reveals a striking physical configuration where QD1 enters an ABS while QD2 occupies a well-defined single-particle state, yet maintains non-trivial quantum coupling with QD1. The observed discontinuities in all measured physical quantities across the phase transition boundary are fundamentally rooted in this profound reconstruction of the ground state wavefunction, demonstrating the critical relationship between quantum state reorganization and macroscopic observable phenomena.
@@
Figure 3(e) demonstrates the Josephson current $I_{J}$ as a function of the Coulomb interaction strength $U_{2}$ for a superconducting phase difference $\phi=0.3\pi$, revealing several key features of the QPT. The critical point shifts to $U_{2}\approx 6\Delta$, demonstrating the $\phi$-dependence of the transition threshold through the relation $U_{2c}=U_{2c}(\phi)$. Prior to the QPT, $I_{J}$ exhibits a continuous sign reversal from positive to negative with increasing $U_{2}$, this behavior emerges from the competing contributions of different ground-state components: the states $|0_{1}0_{2}\rangle$, $|0_{1}d_{2}\rangle$, $|d_{1}0_{2}\rangle$, and $|d_{1}d_{2}\rangle$ contribute to the current through two distinct mechanisms - (i) Cooper pair tunneling producing a positive current and (ii) single-electron tunneling generating a negative current, where initially the Cooper pair contribution dominates, leading to a net positive current. Since only QD1 is directly coupled to both SC leads, the antiferromagnetic state $|\!\!\uparrow_{1}\downarrow_{2}\rangle-|\!\!\downarrow_{1}\uparrow_{2}\rangle$ contributes exclusively through single-electron tunneling. This evolution is quantified by the $U_{2}$-dependent weights $A_{1}$, $B_{1}$, $C_{1}$, $D_{1}\rightarrow 0$ and $E_{1}\rightarrow 1$, causing the net current to reverse sign. Across the QPT, a second current reversal occurs from negative ($I_{J}<0$ for $U_{2}\lesssim 6\Delta$) to positive values ($I_{J}>0$ for $U_{2}\gtrsim 6\Delta$) as QD1 enters an ABS , which strongly enhances Cooper pair transport in the strong correlation regime.

### v2

# Quantum phase transition in a double quantum dot Josephson junction driven by electron-electron interactions

@@ Abstract
In this work, we employ a surrogate BCS model with discrete energy levels to investigate a hybrid system comprising two quantum dots (QD1 and QD2), where QD1 is tunnel-coupled to two superconducting leads, and QD2 is only tunnel-coupled to QD1. Through exact diagonalization of this system, we obtain numerically exact solutions that enable rigorous computation of key physical quantities. Our analysis reveals a rich phase diagram featuring multiple controllable phase transitions mediated by quantum dot electron-electron interaction strength. Specifically, the system first undergoes an initial phase transition when tuning QD2’s electron-electron interaction strength while maintaining QD1 in the non-interacting regime. Subsequent adjustment of QD1’s electron-electron interaction strength induces a secondary phase transition, followed by a third transition arising from inter-dot electron-electron interaction modulation. Furthermore, we demonstrate that parallel magnetic field application can drive reversible ferromagnetic-antiferromagnetic phase transitions under specific parameter conditions. Finally, we report the emergence of non-local magnetization phenomena when subjecting QD1 to weak magnetic fields. And our results demonstrate that the orientation of nonlocal magnetization can be precisely manipulated through adjustment of the on-site interaction strength $U_{2}$ in QD2.
@@ I Introduction
The Josephson junction represents a paradigmatic mesoscopic system comprising a central weak-link region sandwiched between two superconducting electrodes. This fundamental structure has served as a versatile platform for investigating coherent quantum transport phenomena, as evidenced by extensive theoretical [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] and experimental [11, 12, 13, 14, 15, 16]studies. A hallmark characteristic of Josephson junctions is the generation of a dissipationless supercurrent—the Josephson current—that flows without any applied voltage bias when a phase difference exists between the superconducting terminals[17, 18]. This equilibrium current arises from phase-coherent Andreev reflection processes at the superconductor-normal interfaces. A particularly intriguing phenomenon in these systems is the $0-\pi$ quantum phase transition [19, 20, 21, 22, 23, 24, 25, 26, 27, 28], which manifests as a sign reversal of the Josephson current and has been unambiguously observed in experiments [29, 30, 31, 32, 33, 34, 35]. This transition fundamentally reflects a change in the ground state parity of the system. In quantum dot(QD)-based Josephson junctions, two distinct control parameters can drive this transition: (i) tuning the on-dot Coulomb interaction strength [19, 20, 21, 22, 23, 24], or (ii) applying and varying an external magnetic field [25, 26, 27, 28]. Remarkably, despite their different control mechanisms, both approaches share a common microscopic origin—they induce a quantum phase transition(QPT) between two distinct ground state configurations: a singlet state (formed by Cooper pairs via Andreev bound states) and a doublet state (characterized by an unpaired electron). This transition can be understood within a unified framework of competition between Kondo screening or magnetic field and superconducting correlations, where the relative strength of these effects determines the ground state parity.
@@
While QPTs in Josephson junctions containing a single QD in their central region are now well understood, significant research attention has shifted to double-QD systems, as evidenced by growing theoretical [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53] and experimental[54, 55, 56, 57, 58, 59] investigations. The enhanced complexity of ground state physics in these two-QD systems gives rise to QPT phenomena that extend beyond conventional $0-\pi$ transitions, manifesting in three primary configurations: (i) the series configuration [36, 37, 38, 39, 40, 41, 42], where interconnected QDs separately link to each superconducting lead; (ii) the parallel configuration [43, 44, 45, 46, 47, 48], featuring both QDs jointly connected to each superconducting lead; and (iii) the Fano-type configuration [49, 50, 51], where QD1 bridges the superconductors while QD2 remains indirectly coupled. Recent work on Fano-type double interacting QD josephson junction [50] has revealed both a second-order superconducting proximity effect and QPTs induced through QD2’s electron interaction strength tuning, yet crucially omitted examination of QD1’s on-site interactions and the effects of inter-dot interactions between QD1 and QD2 - either of which could drive additional QPTs and represent compelling research avenues. Furthermore, given established magnetic-field-induced QPTs in single-QD junctions, the analogous response in double-QD systems remains an open question of considerable fundamental interest, presenting an important unexplored dimension in hybrid superconducting nanostructure research.
@@
Historically, theoretical analysis of Josephson junctions incorporating interacting QDs posed significant challenges due to the inherent complexity of treating electron correlations. Until recently, the typically effective approaches were Wilson’s numerical renormalization group (NRG) [60, 61, 50, 44] and quantum Monte Carlo (QMC) [62, 63, 64] methods, both of which are computationally demanding. The emergence of the surrogate model technique [65, 66, 67, 68, 69, 70] has provided a simpler yet rigorous alternative to these conventional methods. This innovative approach transforms the continuous leads Hamiltonian into a discrete analog through systematic discretization, enabling exact diagonalization of the complete system Hamiltonian. The resulting eigenvalues and eigenstates provide direct access to all relevant physical quantities through straightforward computation. Importantly, this method has already demonstrated its effectiveness for single-QD Josephson junctions, and its fundamental premise - the faithful representation of superconducting leads through equivalent discrete Hamiltonians - strongly suggests equal validity for double-QD systems. In the present study, we implement this approach by discretizing each superconducting lead into three equivalent sites, then employ state-space expansion techniques to perform exact diagonalization of the full Hamiltonian for a Fano-coupled double-QD Josephson junction. This methodology allows comprehensive computation of system observables, which we subsequently analyze to investigate: (i) interaction-driven QPTs under various quantum dot coupling regimes, and (ii) magnetic-field-induced phase transition phenomena in this novel nanoscale system.
@@ II Model Hamiltonian and Theoretical Method
$\begin{split}H_{QD}=&\sum_{i\sigma}\varepsilon_{i\sigma}d^{{\dagger}}_{i\sigma}d_{i\sigma}+\sum_{i}U_{i}n_{i\uparrow}n_{i\downarrow}\\ &+U_{12}n_{1}n_{2}+t\sum_{\sigma}(d^{{\dagger}}_{1\sigma}d_{2\sigma}+h.c.),\end{split}$ (2)
@@
The operator $d^{{\dagger}}_{i\sigma}$($d_{i\sigma}$) creates (annihilates) an electron with spin $\sigma$ and energy $\varepsilon_{i\sigma}$ on the $i$-th QD, where $U_{i}$ represents the on-site Coulomb repulsion. The interdot tunneling is characterized by the hopping amplitude $t$, while $U_{12}$ denotes the interdot Coulomb interaction, which is typically weaker than the intradot interactions($U_{12}<U_{i}$). Similarly, $c_{\eta\bm{k}\sigma}^{{\dagger}}$($c_{\eta\bm{k}\sigma}$) creates (annihilates) an electron with spin $\sigma$, momentum $\bm{k}$, and energy $\varepsilon_{\eta\bm{k}\sigma}$ in the $\eta$-th SC lead, where $\Delta e^{i\phi_{\eta}}$ represents the complex superconducting order parameter, where $\Delta$ and $\phi$ are real numbers. The SC-QD coupling is described by the tunneling Hamiltonian $H_{T}$ with momentum-independent tunneling matrix elements $V$. We approximate the momentum summation by an energy integral using a constant density of states $\rho_{F}=\frac{1}{2D}$, where $D$ is the half-bandwidth. Throughout this work, we employ natural units with $\hbar$ = $k_{B}$ = $e$ = 1.
@@
the $\tilde{g}$ function is obtained by integrating out an effective superconducting bath that has the same gap $\Delta$ as the original system. This bath consists of three discrete energy levels, located at $\xi_{0}$ and ±$\pm|\xi_{1}|$, which are coupled to the dot via a tunneling matrix element $\tilde{t}_{l}=\sqrt{\gamma_{1}\Gamma}/2$. Since the number of discrete levels is odd, it follows that $\xi_{0}=0$ [70]. The effective bath is thus characterized by the parameters $\{{\gamma_{0}},{\gamma_{1}},\xi_{0},\xi_{1}\}$, which are determined by minimizing the cost function:
@@
Once a satisfactory fit is obtained, the parameters $\{{\gamma_{0}},{\gamma_{1}},\xi_{0},\xi_{1}\}$ define a surrogate model Hamiltonian, representing a discretized version of the original system. This is achieved by replacing the continuous momentum variable $\bm{k}$ with a discrete index $l$, such that $\xi_{\bm{k}}\rightarrow\xi_{l}$ and $t\rightarrow\tilde{t}_{l}$. The resulting Hamiltonian takes the form:
@@ III Results and Discussions
In this work, we numerically solve the low-dimensional surrogate model through exact diagonalization and subsequently calculate the system’s thermodynamic properties including entropy, parity, spin correlations, QD pair term, QD occupation number, and Josephson current, with particular emphasis on the particle-hole symmetric point where $\varepsilon_{i}=-\frac{U_{i}}{2}$. Following standard conventions, we establish $\phi_{L}=\phi$ and $\phi_{R}=0$ for the superconducting phases, with $\phi=0$ as the default value, and maintain the system at zero temperature unless otherwise specified. Furthermore, we only focus on the strong tunneling regime ( $\Gamma$ > $\Delta$), we fix the tunneling rate at $\Gamma=2$ and adopt the superconducting gap $\Delta=1$ as our fundamental energy scale unless specifically stated.
@@ III.1 Single QD josephson junction
$I_{J}=2\frac{\partial E_{S}}{\partial\phi},$ (15)
@@
III.2 QPT are induced by varying the electron-electron interaction parameters
For the subsequent calculations, we fix the interdot tunneling parameter at $t=1$ and examine the system’s physical properties.
@@ III.2 QPT are induced by varying the electron-electron interaction parameters
Fig. 3: (Colour online) System properties analysis: (a) entropy at fixed temperature $T=0.005$, (b) the expectation value of system parity operator $P$ for various magnetic field strengths $B$, (c) spin correlation function $\langle\vec{S}_{1}\!\!\cdot\!\!\vec{S}_{2}\rangle$, (d) normalized superconducting order parameters, $\langle d_{1\downarrow}d_{1\uparrow}\rangle/\Delta$ for QD1 and $\langle d_{2\downarrow}d_{2\uparrow}\rangle/\Delta$ for QD2 and (e) Josephson current $I_{J}/I_{0}$ (where $I_{0}\equiv e\Delta/\hbar$) at fixed superconducting phase difference $\phi=0.3\pi$ as a function of QD2 interaction strength $U_{2}/\Delta$. (f) $I_{J}/I_{0}$ versus phase difference $\phi$ for varying QD2 interaction strengths. The remaining system parameters are fixed at interdot tunneling $t=1$, on-site interaction $U_{1}=0$, and interdot interaction $U_{12}=0$.
@@
Figure 3(a) displays the entropy as a function of QD2 interaction strength $U_{2}$ at fixed temperature $T=0.005$. Throughout Fig. 3, we maintain $U_{1}=0$ for QD1 and $U_{12}=0$ for interdot Coulomb interaction. The entropy exhibits a QPT near $U_{2}\approx 5\Delta$, with three distinct regimes: (i) $S=0$ (singlet state) for $U_{2}<5\Delta$, (ii) $S=\ln 3$ at the critical point, and (iii) $S=\ln 2$ (doublet state) for $U_{2}>5\Delta$. The enhanced entropy value of $\ln 3$ at the quantum critical point, as previously discussed, arises from the emergent threefold degeneracy in the ground state manifold during the phase transition. Figure 3(b) shows the parity $P$ versus $U_{2}$. Post-transition, the system enters a doublet state. To enable parity measurement, we apply a weak magnetic field to QD1, lifting the degeneracy. The parity transitions sharply from 00 (even) to ±$\pm 1$ (odd) at the QPT, confirming the change in ground state. The interdot spin correlation in Fig. 3(c) plots an antiferromagnetic configuration ($\langle\vec{S}_{1}\!\cdot\!\vec{S}_{2}\rangle<0$) before the QPT that vanishes ($\langle\vec{S}_{1}\!\cdot\!\vec{S}_{2}\rangle\approx 0$) in the doublet phase. The spin correlation behavior emerges from distinct ground state configurations in different phases: (i) In the singlet phase prior to the QPT, the formation of spatially localized Cooper pairs between QD1 and QD2 yields antiferromagnetic spin alignment; (ii) In the post-transition doublet phase where QD1 enters an ABS, the spin correlations become negligible. Illustrated in Fig. 3(d) is the normalized superconducting order parameters $\langle d_{1\downarrow}d_{1\uparrow}\rangle/\Delta$ (QD1) and $\langle d_{2\downarrow}d_{2\uparrow}\rangle/\Delta$ (QD2) versus $U_{2}$. At weak interactions, both orders are significant, but they decrease with increasing $U_{2}$ due to Coulomb suppression of double occupancy and empty occupancy. After the QPT, $\langle d_{1\downarrow}d_{1\uparrow}\rangle$ surges while $\langle d_{2\downarrow}d_{2\uparrow}\rangle$ vanishes, reflecting the system’s transition to a state where QD2 reaches single occupancy (particle-hole symmetry point) while QD1 forms ABS.
@@
Through rigorous analysis of the aforementioned physical quantities, we establish that the system’s ground state prior to the QPT or in the absence of electronic correlations within the QDs can be precisely described by the superposition state: $A_{1}|0_{1}0_{2}\rangle+B_{1}|0_{1}d_{2}\rangle+C_{1}|d_{1}0_{2}\rangle+D_{1}|d_{1}d_{2}\rangle+E_{1}(|\!\!\uparrow_{1}\downarrow_{2}\rangle-|\!\!\downarrow_{1}\uparrow_{2}\rangle)$, where $d_{1}$, $d_{2}$ represent the double occupation state of QD, the coefficients $A_{1}$, $B_{1}$, $C_{1}$, $D_{1}$, and $E_{1}$ quantitatively characterize the relative contributions of each entangled quantum state. Remarkably, the system undergoes a dramatic reorganization during the QPT, with the ground state abruptly collapsing into the simplified forms $|0_{1}\!\!\uparrow_{2}\rangle-|d_{1}\!\!\uparrow_{2}\rangle$ and $|0_{1}\!\!\downarrow_{2}\rangle-|d_{1}\!\!\downarrow_{2}\rangle$. This transition reveals a striking physical configuration where QD1 enters an ABS while QD2 occupies a well-defined single-particle state, yet maintains non-trivial quantum coupling with QD1. The observed discontinuities in all measured physical quantities across the phase transition boundary are fundamentally rooted in this profound reconstruction of the ground state wavefunction, demonstrating the critical relationship between quantum state reorganization and macroscopic observable phenomena.
@@
Figure 3(e) demonstrates the Josephson current $I_{J}$ as a function of the Coulomb interaction strength $U_{2}$ for a superconducting phase difference $\phi=0.3\pi$, revealing several key features of the QPT. The critical point shifts to $U_{2}\approx 6\Delta$, demonstrating the $\phi$-dependence of the transition threshold through the relation $U_{2c}=U_{2c}(\phi)$. Prior to the QPT, $I_{J}$ exhibits a continuous sign reversal from positive to negative with increasing $U_{2}$, this behavior emerges from the competing contributions of different ground-state components: the states $|0_{1}0_{2}\rangle$, $|0_{1}d_{2}\rangle$, $|d_{1}0_{2}\rangle$, and $|d_{1}d_{2}\rangle$ contribute to the current through two distinct mechanisms - (i) Cooper pair tunneling producing a positive current and (ii) single-electron tunneling generating a negative current, where initially the Cooper pair contribution dominates, leading to a net positive current. Since only QD1 is directly coupled to both SC leads, the antiferromagnetic state $|\!\!\uparrow_{1}\downarrow_{2}\rangle-|\!\!\downarrow_{1}\uparrow_{2}\rangle$ contributes exclusively through single-electron tunneling. This evolution is quantified by the $U_{2}$-dependent weights $A_{1}$, $B_{1}$, $C_{1}$, $D_{1}\rightarrow 0$ and $E_{1}\rightarrow 1$, causing the net current to reverse sign. Across the QPT, a second current reversal occurs from negative ($I_{J}<0$ for $U_{2}\lesssim 6\Delta$) to positive values ($I_{J}>0$ for $U_{2}\gtrsim 6\Delta$) as QD1 enters an ABS region , which strongly enhances Cooper pair transport in the strong correlation regime.